OpenStudy (anonymous):

find the derivative f(x)=log3x

OpenStudy (anonymous):

\[f(x)=\log _{3}x\]

OpenStudy (anonymous):

I'm beginning to think that you're a math robot

jimthompson5910 (jim_thompson5910):

bleep blorp...i mean...crap I've been discovered lol

OpenStudy (anonymous):

Seriously!

jimthompson5910 (jim_thompson5910):

Use the rule that If \[\Large f(x) = \log_{b}(x)\] then \[\Large f^{\prime}(x) = \frac{1}{x*\ln(b)}\]

jimthompson5910 (jim_thompson5910):

So in this case, b = 3

jimthompson5910 (jim_thompson5910):

again this is just another rule to learn...but it's very similar to the rule that If \[\Large f(x) = \ln(x)\] then \[\Large f^{\prime}(x) = \frac{1}{x}\]

OpenStudy (anonymous):

okay so 1/xln3

jimthompson5910 (jim_thompson5910):

so that's something to connect

OpenStudy (anonymous):

so x*1/3?

OpenStudy (anonymous):

x/3

jimthompson5910 (jim_thompson5910):

yes, in this case, \[\Large f^{\prime}(x) = \frac{1}{x*\ln(3)}\]

jimthompson5910 (jim_thompson5910):

no you can't get rid of the ln that easily

OpenStudy (anonymous):

dang!

jimthompson5910 (jim_thompson5910):

yeah I wish you could, but you can't

terenzreignz (terenzreignz):

maybe it's prudent to show where that came from? After all... \[\huge \log_3x=\frac{\ln \ x}{\ln \ 3}\] And now, the \[\huge \frac1{\ln 3}\]is just another constant.

jimthompson5910 (jim_thompson5910):

good point terenzreignz, that shows how the two connect

OpenStudy (anonymous):

alright, so this boils down to the derivative being 1/xln3

OpenStudy (anonymous):

I cant simplify or anything?

jimthompson5910 (jim_thompson5910):

yes \[\Large f^{\prime}(x) = \frac{1}{x*\ln(3)}\]

terenzreignz (terenzreignz):

^ That's as simple as it can be :)

OpenStudy (anonymous):

first use change of base then use natural log log3x can be written as using change of base property of log \[\Large \log_{3}(x)=\frac{\ln(x)}{\ln(3)}\] now take the constant ln(3) out \[\Large \frac{d}{dx}(\frac{\ln(x)}{\ln(3)})=\frac{1}{\ln(3)}*\frac{d}{dx}(\ln(x)\] \[\Large \frac{1}{\ln(3)}*\frac{1}{x}=\frac{1}{\ln(3)x}\]

jimthompson5910 (jim_thompson5910):

you can rearrange terms I guess, but it's not going to get any simpler

OpenStudy (anonymous):

alright...that seems okay

OpenStudy (anonymous):

not too hard, until i look at the next set and theres RADICALS!

terenzreignz (terenzreignz):

Radicals are just fractional exponents... just sayin' ;)

OpenStudy (anonymous):

\[f(x)=2^{x}\sqrt{x}\]

OpenStudy (anonymous):

so that changes to 1/x?

terenzreignz (terenzreignz):

Use the product rule, but remember that \[\huge f(x)=2^x \cdot x^{\frac12}\]

OpenStudy (anonymous):

how does that apply to the fraction though

OpenStudy (anonymous):

the radical i mean

terenzreignz (terenzreignz):

Power rule :) \[\huge \frac{d}{dx}x^m=mx^{m-1}\]for any real number m.

OpenStudy (anonymous):

So I can use that for the 2^x and make it 2x but I still have that times the radical x

terenzreignz (terenzreignz):

No! The power rule applies when x is the BASE, not when x is the EXPONENT. We'll get to that later. For now, what's the derivative of the square root of x? \[\huge \frac{d}{dx}\sqrt x=\frac{d}{dx}x^{\frac12}=?\]

OpenStudy (anonymous):

1/x?

terenzreignz (terenzreignz):

Nope...Remember the power rule...\[\huge \frac{d}{dx}x^m=mx^{m-1}\]and take m = 1/2

OpenStudy (anonymous):

1/2(radical x)^-1/2

terenzreignz (terenzreignz):

This? \[\huge \frac12 \sqrt x^{-\frac12}\]?

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

Well, in that case, no. We already converted the radical to a fractional exponent, why did you put it back? :P

OpenStudy (anonymous):

I guess I missed that part!

terenzreignz (terenzreignz):

So... care to modify your answer? :P

terenzreignz (terenzreignz):

You were *almost* right, though.

OpenStudy (anonymous):

so just take out the radical

terenzreignz (terenzreignz):

Yeah :D \[\huge \frac12x^{-\frac12}=\frac1{2\sqrt x}\] Ok? :P

OpenStudy (anonymous):

okay!

terenzreignz (terenzreignz):

So, no questions on the part of... \[\huge \frac{d}{dx}\sqrt x = \frac{1}{2 \sqrt{x}} \]

OpenStudy (anonymous):

well now that i am able to see where you got the 1/2 from I understand

terenzreignz (terenzreignz):

So, shall we proceed? :)

OpenStudy (anonymous):

absolutely

terenzreignz (terenzreignz):

Now, work on the derivative of \[\huge 2^x\]Any idea how to proceed?